OpenStudy (anonymous):

Matrix equation B^101 ?

OpenStudy (dogzcatz):

is it multiple choice

OpenStudy (anonymous):

\[B=\left[\begin{matrix}-1 & -\sqrt3 \\ \sqrt3 & -1\end{matrix}\right]\]

OpenStudy (anonymous):

I know that B^3 gives the unit matrix and that B^4 = B, right?

OpenStudy (anonymous):

But how to calculate B^101?

OpenStudy (anonymous):

sorry \[B =\frac{ 1 }{ 2 }\left[\begin{matrix}-1 & -\sqrt3 \\ \sqrt3 & -1\end{matrix}\right]\] B is equal to that

OpenStudy (anonymous):

And from Matlab I know that \[B ^{101} = \frac{ 1 }{ 2 }\left[\begin{matrix}-1 & \sqrt3 \\ -\sqrt3 & -1\end{matrix}\right]\]